### v1

# Heapr: Hessian-based Efficient atomic expert Pruning in Output Space

Abstract
Mixture-of-Experts (MoE) architectures in large language models (LLMs) deliver exceptional performance and reduced inference costs compared to dense LLMs. However, their large parameter counts result in prohibitive memory requirements, limiting practical deployment. While existing pruning methods primarily focus on expert-level pruning, this coarse granularity often leads to substantial accuracy degradation. In this work, we introduce Heapr, a novel pruning algorithm that decomposes experts into smaller, indivisible atomic experts, enabling more precise and flexible atomic expert pruning. To measure the importance of each atomic expert, we leverage second-order information based on principles similar to Optimal Brain Surgeon (OBS) theory. To address the computational and storage challenges posed by second-order information, Heapr exploits the inherent properties of atomic experts to transform the second-order information from expert parameters into that of atomic expert parameters, and further simplifies it to the second-order information of atomic expert outputs. This approach reduces the space complexity from $O(d^{4})$, where $d$ is the model’s dimensionality, to $O(d^{2})$. Heapr requires only two forward passes and one backward pass on a small calibration set to compute the importance of atomic experts. Extensive experiments on MoE models, including DeepSeek MoE and Qwen MoE family, demonstrate that Heapr outperforms existing expert-level pruning methods across a wide range of compression ratios and benchmarks. Specifically, Heapr achieves nearly lossless compression at compression ratios of $20\%\sim 25\%$ in most models, while also reducing FLOPs nearly by $20\%$. The code can be found at https://github.com/LLIKKE/HEAPr.
1 Introduction
Mixture-of-experts (MoE) models have recently emerged as a promising alternative to dense large language models (LLMs), replacing dense feed-forward layers with sparsely activated experts and dynamic routing. This design allows MoE models to match or surpass the performance of dense LLMs while activating only a fraction of parameters during inference (Fedus et al., 2022; Zhu et al., 2024; Liu et al., 2024), making them particularly attractive for large-scale, concurrent deployment. However, while sparse activation reduces computational cost, it exacerbates memory requirements. For example, DeepSeek-V3 (Liu et al., 2024) activates only 37B parameters per inference, yet all 671B parameters must still be stored in GPU memory, resulting in prohibitively high deployment costs. Notably, MoE layers typically account for over $97\%$ of total model parameters, they represent the dominant storage bottleneck. Therefore, compressing MoE layers becomes critical to overcoming inference inefficiency and making deployment feasible in resource-constrained environments.
Model pruning has been widely explored as an effective compression strategy to reduce storage and improve efficiency. Yet a fundamental trade-off persists: fine-grained pruning typically preserves accuracy but yields limited speedups on hardware, whereas coarse-grained pruning directly enables acceleration but often incurs obvious accuracy loss. Within MoE models, parameter sparsification (Xie et al., 2024) faces similar limitations, as hardware inefficiencies constrain its practical
Figure 1: Illustration of atomic expert-level pruning, which removing the $t$-th column from the $W_{gate}$ and $W_{up}$ matrices, and the corresponding $t$-th row from the $W_{down}$ matrix.
benefits. Consequently, recent research has shifted toward expert-level pruning, offering more direct gains in both acceleration and memory reduction. Existing expert-level approaches at this level can be broadly divided into expert dropping and expert merging. Expert dropping methods (Lu et al., 2024; Huang et al., 2025) completely remove experts deemed unimportant, but relying solely on calibration to discard entire experts risks losing valuable complementary expertise, consequently often leading to notable performance degradation. Expert merging methods (Li et al., 2024; Chen et al., 2025; Huang et al., 2025) instead aim to consolidate functionally similar experts to more effectively preserve overall model capacity. However, their clustering-based similarity measures are notoriously unstable, and naive merging strategies (e.g., averaging or frequency-based weighting) often introduce destructive parameter conflicts, resulting in suboptimal and inefficient outcomes. To alleviate these critical conflicts, recent decomposition-based approaches (Li et al., 2025b; Gu et al., 2025) represent individual experts as a mixture of shared and specialized components. While this advanced framework helps to preserve model capacity, it still requires computationally expensive decomposition and merging operations, and unfortunately still incurs a non-negligible accuracy loss.
To identify pruning units that are more flexible than expert-level pruning, we introduce the concept of an atomic expert, in which each expert is decomposed into smaller, indivisible units. Concretely, each atomic expert is defined by jointly grouping the relevant columns of $\bm{W}^{\mathrm{up}}$, $\bm{W}^{\mathrm{gate}}$, and the corresponding row of $\bm{W}^{\mathrm{down}}$(as shown in Figure 1). The output of a full expert can be represented as the sum of outputs from multiple atomic experts. Pruning at this granularity directly removes atomic experts, thereby isolating pruning effects and avoiding interference with remaining components. By eliminating atomic experts that contribute little to final predictions, inference efficiency can be improved and deployment overhead reduced in a more straightforward and essential way.
[…]
  • We introduce a second-order approximation scheme for atomic expert pruning in MoE models, which transforms the second-order information from expert parameters into that of atomic expert parameters, and further simplifies it to the second-order information of atomic expert outputs. This approach reduces the space complexity of second-order information from $\mathcal{O}\left((3d_{\text{model}}\cdot d_{\text{inter}})^{2}\right)$ to $\mathcal{O}(d_{\text{model}}^{2})$.
  • Building on this efficient scheme, we propose Heapr, a highly efficient and scalable pruning algorithm that accurately estimates the importance of all atomic experts with just two forward passes and one backward pass on a small calibration set.
  • We conduct extensive experiments on DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-Chat, Qwen2-57B-A14B, and Qwen3-30B-A3B across diverse benchmarks. Heapr outperforms current SOTA methods and achieves nearly lossless compression at compression ratios of $20\%–25\%$ in most models, while also reducing FLOPs by nearly $20\%$.
2 related works and Preliminary
Mixture of Experts Compression.
Model compression for MoE architectures has recently attracted growing attention due to the remarkable performance of MoE models. MoE-Pruner (Xie et al., 2024) performs weights sparsification based on activation magnitude, weight magnitude, and router importance, yet its acceleration is hardware-dependent and relies on distillation to recover accuracy. Expert-level pruning has been more extensively explored due to its hardware-friendly acceleration. NAEE (Lu et al., 2024) selects a subset of experts to minimize calibration error, but this can lead to overfitting and the loss of specialized knowledge. Similarly, MoE-I2 (Yang et al., 2024) combines expert pruning with low-rank decomposition, yet requires additional fine-tuning for recovery. To alleviate such issues, expert merging methods aim to retain similar experts rather than discarding them. MC-SMoE (Li et al., 2024) merges experts by clustering based on routing policies, and HC-MoE (Chen et al., 2025) does so by grouping experts with similar outputs. However, limited expert similarity makes merging prone to parameter conflicts. To further exploit redundancy, $D^{2}$-MoE (Gu et al., 2025) constructs a shared expert via weighted combinations and compresses residuals through low-rank decomposition, while Sub-MoE (Li et al., 2025a) applies SVD to extract a shared subspace across experts, both of which require computationally expensive decomposition and merging operations. We decompose the expert into atomic experts and propose Heapr, a method that measures atomic expert importance by utilizing a second-order approximation to assess the importance of atomic experts. This approach enables more flexible pruning units and provides a highly efficient algorithmic, preserving model performance while eliminating the need for retraining.
Optimal Brain Surgeon in pruning.
The OBS framework (Hassibi et al., 1993; LeCun et al., 1989) approaches pruning as an optimization problem, aiming to minimize the increase in the loss function when a parameter is removed. Consider a model that has already been trained and converged, with parameters $\theta$ and a corresponding loss $\ell(\theta)$. We can analyze the effect of perturbing the parameters by analyzing the second-order Taylor expansion of the loss function around $\theta$. Specifically, the change in the loss $\Delta\ell$ when perturbing the parameters by $\delta\mathbf{\theta}$ is given by the following:
$\Delta\ell=\ell(\theta+\delta\theta)-\ell(\theta)=\nabla\ell(\theta)^{\top}\delta\theta+\frac{1}{2}\delta\theta^{\top}\bm{H}\delta\theta+O(\|\delta\theta\|^{3}),$ (1)
where $\bm{H}$ is the Hessian matrix of second derivatives of the loss with respect to the model parameters. Since the model has already converged to a local minimum of the loss function, the first-order term can be removed ($\nabla\ell(\theta)=\mathbf{0}$), and the higher-order terms can be ignored for small perturbations.
For pruning, the constraint is $\theta_{q}+\delta\theta_{q}=0$ for the target, leading to the optimization problem as:
$\min_{\delta\theta_{q}}\;\;\tfrac{1}{2}\delta\theta^{\top}\bm{H}\delta\theta,\quad\text{s.t. }\delta\theta_{q}+\theta_{q}=0,$ (2)
where $q$ denotes the index of the pruned parameter. Solving this optimization problem yields the minimal increase in loss from pruning parameter $\theta_{q}$, which is $\Delta\ell=\tfrac{1}{2}\frac{\theta_{q}^{2}}{[\bm{H}^{-1}]_{qq}}.$
Directly computing the full Hessian in deep neural networks is practically infeasible. Existing OBS methods to adopt significant approximations. For instance, K-FAC approximation (Martens & Grosse, 2015) provides an efficient approximation of second-order information and Hessians are computed layer-wise to guide pruning (Dong et al., 2017; Frantar & Alistarh, 2022; Frantar et al., 2023). Previous work (Singh & Alistarh, 2020) shows that the Fisher information matrix serves as a reliable Hessian estimate and allows for more efficient computation. Some apply OBS to structured pruning (Yu et al., 2022), but these efforts are limited as they consider only the trace of the Hessian.
3 Method
3.1 Atomic Expert in Mixture-of-Experts.
The MoE architecture has been widely adopted in LLMs as a replacement for the dense feed-forward network layer, which effectively increases the model capacity while reducing the number of activated parameters. Formally, given an input token representation $\mathbf{x}\in\mathbb{R}^{d_{\text{model}}}$, the MoE output is defined as:
$\mathbf{y}=\sum_{i=1}^{\kappa}g_{i}(\mathbf{x})\,E_{i}(\mathbf{x}),\qquad\mathbf{g}(\mathbf{x})=\big(g_{1}(\mathbf{x}),\dots,g_{\kappa}(\mathbf{x})\big)=\operatorname{Top}\text{-}\kappa\!\big(\bm{W}^{\mathrm{gate}}\mathbf{x}\big)\in\mathbb{R}^{\kappa},$ (3)
where $\bm{W}^{\mathrm{gate}}\in\mathbb{R}^{E\times d_{\text{model}}}$ produces router scores and $\operatorname{Top}\text{-}\kappa(\cdot)$ denotes the router function that selects the top-$\kappa$ experts. Each expert $E_{i}(\cdot)$ is a gated feed-forward block:
$E_{i}(\mathbf{x})=\bm{W}_{i}^{\mathrm{down}}\big[\,\operatorname{SiLU}\!\big(\bm{W}_{i}^{\mathrm{gate}}\mathbf{x}\big)\odot\big(\bm{W}_{i}^{\mathrm{up}}\mathbf{x}\big)\big],$ (4)
where $\bm{W}_{i}^{\mathrm{up}},\bm{W}_{i}^{\mathrm{gate}}\in\mathbb{R}^{d_{\text{inter}}\times d_{\text{model}}}$, $\bm{W}_{i}^{\mathrm{down}}\in\mathbb{R}^{d_{\text{model}}\times d_{\text{inter}}}$, $\odot$ denotes the Hadamard product, and $\operatorname{SiLU}(\cdot)$ is the SiLU activation. Within each expert, computations can be decomposed into atomic experts. Let $\mathbf{w}_{i,j}^{\mathrm{up}}$ and $\mathbf{w}_{i,j}^{\mathrm{gate}}$ denote the $j$-th rows of $\bm{W}_{i}^{\mathrm{up}}$ and $\bm{W}_{i}^{\mathrm{gate}}$, respectively, and let $\mathbf{w}_{i,j}^{\mathrm{down}}$ denote the $j$-th column of $\bm{W}_{i}^{\mathrm{down}}$. Then the $j$-th atomic expert of the $i$-th expert is
$\mathbf{e}_{i}^{(j)}(\mathbf{x})=\mathbf{w}_{i,j}^{\mathrm{down}}\,\Big[\operatorname{SiLU}\!\big(\mathbf{w}_{i,j}^{\mathrm{gate}}\mathbf{x}\big)\cdot\big(\mathbf{w}_{i,j}^{\mathrm{up}}\mathbf{x}\big)\Big]\;\in\;\mathbb{R}^{d_{\text{model}}},$ (5)
where $\mathbf{w}_{i,j}^{\mathrm{up}},\mathbf{w}_{i,j}^{\mathrm{gate}}\in\mathbb{R}^{1\times d_{\text{model}}}$ and $\mathbf{w}_{i,j}^{\mathrm{down}}\in\mathbb{R}^{d_{\text{model}}\times 1}$. Consequently, each expert is a linear combination of its atomic experts:
$E_{i}(\mathbf{x})=\sum_{j=1}^{d_{\text{inter}}}\mathbf{e}_{i}^{(j)}(\mathbf{x}).$ (6)
In this framework, each expert $E_{i}(\cdot)$ can be viewed as a linear combination of its atomic experts. This decomposition allows pruning at the atomic expert level without compromising the other atomic expert structure, leading to both computational acceleration and deployment efficiency directly.
3.2 Atomic expert Importance Analysis in the Output Space
Importance of Atomic experts.
As discussed in Section 2, the OBS theory provides an excellent framework for analyzing the impact of parameter pruning on model performance. However, its major limitation is the large Hessian matrix, even when only computed layer-wise. In the case of MoE, directly applying OBS at the expert level is still infeasible, as it requires constructing an exceedingly large Hessian with space complexity of $\mathcal{O}\left((3d_{\text{model}}\cdot d_{\text{inter}})^{2}\right)$ per expert, leading to prohibitive computation and storage costs. Fortunately, by decomposing the expert into smaller atomic experts, a property is revealed: the parameters of different atomic experts are decoupled, i.e.,
$\frac{\partial^{2}E(\mathbf{x})}{\partial\Theta^{(i)}\,\partial\Theta^{(j)}}=\frac{\partial^{2}\mathbf{e}^{(i)}(\mathbf{x})}{\partial\Theta^{(i)}\,\partial\Theta^{(j)}}=0,\quad\forall i\neq j$ (7)
where $\Theta^{(i)}\in\mathbb{R}^{3d_{\text{model}}}$ represents the parameters of the $i$-th atomic expert. This means that the cross-Hessians between different atomic experts are zero, which provides a valuable and simplifying property that allows us to focus exclusively on the Hessian of each individual atomic expert with respect to its own specific parameters. Based on this observation, the second-order Taylor expansion of the change in the loss function with respect to each expert’s parameters can be expressed as:
$\Delta\ell=\frac{1}{2}\delta\Theta^{T}\bm{H}\delta\Theta=\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}$ (8)
here, $\Theta\in\mathbb{R}^{3d_{\text{model}}\cdot d_{\text{inter}}}$ denotes the parameters of a given expert, while $\bm{H}$ is the corresponding Hessian matrix, which has a space complexity of $\mathcal{O}\left((3d_{\text{model}}\cdot d_{\text{inter}})^{2}\right)$. The matrix $\bm{H}^{(i)}$ represents the Hessian for the $i$-th atomic expert. This decomposition leads to a significant reduction in the complexity of summing over the Hessians $\sum_{i=1}^{d_{\text{inter}}}\bm{H}^{(i)}$, which is reduced to $\mathcal{O}\left((3d_{\text{model}})^{2}\cdot d_{\text{inter}}\right)$. However, the resulting Hessian matrix computation remains unacceptable due to its high computational and storage cost. To further alleviate the bottleneck, we introduce a second optimization that shifts pruning constraints from the parameter space of atomic experts to their output space, where the constraint in equation 2 can be equivalently enforced in the atomic expert output domain. Specifically, pruning the atomic expert $\mathbf{e}_{\mathcal{P}}$ is equivalent to imposing: $\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})=\mathbf{0},$ where $\Theta_{\mathcal{P}}\in\mathbb{R}^{3d_{\text{model}}}$ denotes the parameters of the pruned atomic expert. Since the atomic expert functions are not optimized with respect to the parameters $\Theta_{\mathcal{P}}$ through gradient descent, applying a Taylor expansion of the atomic expert functions around $\Theta_{\mathcal{P}}$ results in the first-order term dominating, yielding:
$\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}}+\delta\Theta_{\mathcal{P}})\approx\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}})+\bm{J}_{\mathcal{P}}\delta\Theta_{\mathcal{P}}=\mathbf{0},$ (9)
where $\bm{J}_{\mathcal{P}}\in\mathbb{R}^{d_{\text{model}}\times 3d_{\text{model}}}$ denotes the Jacobian of $\mathbf{e}_{\mathcal{P}}(\Theta_{\mathcal{P}})$. This leads to the following problem:
$\min_{\Theta_{\mathcal{P}}}\frac{1}{2}\sum_{i=1}^{d_{\text{inter}}}(\delta\Theta^{(i)})^{T}\bm{H}^{(i)}\delta\Theta^{(i)}\quad\text{s.t.}\quad\bm{J}_{\mathcal{P}}\,\delta\Theta_{\mathcal{P}}+\mathbf{e}_{\mathcal{P}}=\mathbf{0}.$ (10)
To solve the problem in equation 10 , we consider the LLMs trained with a negative log-likelihood loss $\ell$ (e.g., cross-entropy loss). In this setting, the Fisher Information Matrix $\bm{F}$ is equivalent to the expected Hessian (Bishop & Nasrabadi, 2006), providing a computationally efficient alternative:
$\mathbb{E}\left[{\bm{H}}\right]=\bm{F}=\mathbb{E}\!\left[(\nabla_{\Theta}\ell)(\nabla_{\Theta}\ell)^{T}\right],$ (11)
where $\ell$ is the sample-wise loss. Previous work (Singh & Alistarh, 2020) has shown that for well-converged neural networks, a few hundred representative samples are already sufficiently reliable to estimate $\mathbb{E}\left[{\bm{H}}\right]$. Expanding the gradient of $\ell$ with respect to the parameters gives $\nabla_{\Theta_{\mathcal{P}}}\ell=\bm{J}_{\mathcal{P}}^{\top}\mathbf{g}_{\mathcal{P}},$ where $\mathbf{g}_{\mathcal{P}}\in\mathbb{R}^{d{\text{model}}}$ is the gradient of the loss with respect to the pruned atomic expert output $\mathbf{e}_{\mathcal{P}}$. Substituting this expression into the objective equation 10 yields the expected loss increase when pruning the atomic expert $\mathbf{e}_{\mathcal{P}}$, with $\delta\Theta^{(i)}=\mathbf{0}$ for all atomic experts not pruned:
$\Delta\ell=\tfrac{1}{2}\,\delta\Theta_{\mathcal{P}}^{T}\mathbb{E}\left[{\bm{H}}_{\mathcal{P}}\right]\,\delta\Theta_{\mathcal{P}}=\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}}.$ (12)
This leads us to define the Importance of the atomic expert $\mathbf{e}_{\mathcal{P}}$ as
$s=\tfrac{1}{2}\,\mathbf{e}_{\mathcal{P}}^{\top}\,\mathbb{E}[\mathbf{g}_{\mathcal{P}}\mathbf{g}_{\mathcal{P}}^{\top}]\,\mathbf{e}_{\mathcal{P}},$ (13)
where a smaller $s$ indicates that the corresponding atomic expert has less impact on the overall model loss and should be pruned with higher priority. The detailed derivation is provided in Appendix A.
At this point, we have shifted the analysis from the parameter space of atomic experts to their output space, further reducing both computational and storage requirements. Next, we introduce a remarkable property of atomic expert outputs: the outputs of atomic experts within the same expert share identical gradients, i.e.,
$\frac{\partial\ell}{\partial\mathbf{e}^{(i)}(\mathbf{x})}=\frac{\partial\ell}{\partial E(\mathbf{x})},\quad\forall i\in\{1,\dots,d_{\text{inter}}\},\ \mathbf{e}^{(i)}\in E.$ (14)
This property allows us to further significantly reduce storage requirements. Instead of maintaining separate gradient covariance matrices for each atomic expert, we only need to store a single matrix per expert. As a result, the space complexity for computing the importance of the atomic expert within the same expert is drastically reduced to $\mathcal{O}(d_{\text{model}}^{2})$, enabling efficient storage management.
Global Ranking of Atomic Experts.
The metric of each atomic expert’s importance has been introduced by equation 13. Next, an important question is how to rank the importance of the atomic experts. Consider that our importance metric evaluates experts based on their overall contribution to the model’s change in the loss function (as shown in equation 2), it provides a natural basis for global ranking. This allows us to effectively compare experts across layers consistently, ensuring that pruning decisions are made based on the entire model’s behavior rather than isolated layer-wise.
3.3 Heapr Algorithm
Building on the above analysis, we propose Heapr, a pruning strategy for MoE feedforward layers that ranks the importance of atomic experts (as defined in equation 13) and removes those with negligible contribution to the overall loss. To effectively compute the importance of atomic experts, we leverage a small but representative calibration set $\mathcal{D}$ and estimate importance in two stages.
1. Shared Gradient Covariance Estimation. For a given expert $E_{i}$, the gradients of the loss with respect to all its constituent atomic experts’ output are identical. Therefore, rather than performing redundant computations, we execute a single backward pass to obtain the gradient for the expert’s output, $\mathbf{g}_{E_{i}}=\partial\ell/\partial E_{i}$. This shared gradient is used to compute a gradient covariance matrix $\bar{\bm{G}}_{i}$, for all atomic experts belonging to $E_{i}$, accumulated over the subset of tokens $\mathcal{T}_{i}\subseteq\mathcal{D}$ routed to $E_{i}$:
$\bar{\bm{G}}_{i}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\mathbf{g}_{E_{i}}(\mathbf{x})\mathbf{g}_{E_{i}}(\mathbf{x})^{\top}.$ (15)
2. Importance Computation. Subsequently, during a forward pass, we compute the importance for each individual atomic expert $\mathbf{e}_{k}$. Although the gradient covariance matrix $\bar{\bm{G}}_{i}$ is shared among all atomic experts within same expert $E_{i}$, the output of each atomic expert, $\mathbf{e}_{k}(\mathbf{x})$, remains unique. This difference in output allows us to distinguish their individual contributions. The importance of an atomic expert $\mathbf{e}_{k}$ (where $\mathbf{e}_{k}\in E_{i}$) is calculated by averaging over the tokens it processes:
$\bar{s}_{k}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\tfrac{1}{2}\,\mathbf{e}_{k}(\mathbf{x})^{\top}\bar{\bm{G}}_{i}\mathbf{e}_{k}(\mathbf{x}).$ (16)
This approach relies solely on standard forward and backward computations, making it both exceptionally time- and memory-efficient. The space complexity of each gradient covariance matrix is only $\mathcal{O}(d_{\text{model}}^{2})$, significantly alleviating the storage bottleneck. After computing the importance $\bar{s}_{k}$} across all micro-experts in the model, we perform a global ranking and prune the lowest $r\%$ of experts across all MoE layers. The complete and optimized procedure is summarized in Algorithm 1.
Algorithm 1 Heapr: Hessian-based Efficient Atomic Expert Pruning
1:MoE model $f_{\theta}$, calibration set $\mathcal{D}$, pruning ratio $r$
2:Pruned model $f_{\theta^{\prime}}$
3:for each expert $E_{i}$ do $\triangleright$ Stage 1: Gradient Covariance Estimation
4:  Collect routed tokens $\mathcal{T}_{i}$
5:  Compute shared gradient $\mathbf{g}_{E_{i}}=\frac{\partial\ell}{\partial E_{i}}$
6:  Compute $\bar{\bm{G}}_{i}=\tfrac{1}{|\mathcal{T}_{i}|}\sum_{\mathbf{x}\in\mathcal{T}_{i}}\mathbf{g}_{E_{i}}(\mathbf{x})\mathbf{g}_{E_{i}}(\mathbf{x})^{\top}$
7:end for
8:for each atomic expert $\mathbf{e}_{k}$ in $E_{i}$ do $\triangleright$ Stage 2: Importance Computation
9:  Compute $\bar{s}_{k}=\tfrac{1}{|\mathcal{T}_{k}|}\sum_{\mathbf{x}\in\mathcal{T}_{k}}\tfrac{1}{2}\,\mathbf{e}_{k}(\mathbf{x})^{\top}\bar{\bm{G}}_{i}\mathbf{e}_{k}(\mathbf{x})$
10:end for
11:Global rank $\{\bar{s}_{k}\}$ and prune lowest $r\%$ across all experts
12:return Pruned model $f_{\theta^{\prime}}$
4 experiment
Table 1: Performance of Heapr with DeepSeekMoE-16B-Base, Qwen1.5-MoE-A2.7B-chat, Qwen2-57B-A14B and Qwen3-30B-A3B on seven zero-shot tasks, reported in terms of accuracy.
Ratio Method Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
DeepSeekMoE-16B-Base
0% Original 0.32 0.76 0.71 0.58 0.45 0.79 0.32 0.56
20% NAEE 0.32 0.71 0.66 0.55 0.40 0.77 0.29 0.53
MoE-I2 0.26 0.71 0.68 0.49 0.38 0.73 0.29 0.50
$D^{2}$-MoE 0.30 0.74 0.69 0.55 0.41 0.76 0.31 0.54
Heapr 0.32 0.76 0.71 0.57 0.45 0.79 0.32 0.56
40% NAEE 0.23 0.67 0.67 0.41 0.32 0.69 0.26 0.46
MoE-I2 0.23 0.64 0.66 0.41 0.31 0.68 0.26 0.45
$D^{2}$-MoE 0.26 0.69 0.65 0.45 0.36 0.72 0.28 0.49
Heapr 0.30 0.74 0.69 0.52 0.41 0.76 0.30 0.53
Qwen1.5-MoE-A2.7B-Chat
0% Original 0.31 0.70 0.66 0.59 0.40 0.79 0.35 0.54
25% MC-SMoE 0.25 0.65 0.65 0.53 0.37 - - -
HC-SMoE 0.27 0.66 0.63 0.55 0.35 0.76 0.29 0.50
Sub-MoE 0.30 0.69 0.66 0.56 0.37 - - -
Heapr 0.32 0.69 0.67 0.56 0.38 0.76 0.35 0.53
50% MC-SMoE 0.18 0.33 0.52 0.29 0.19 - - -
HC-SMoE 0.23 0.61 0.65 0.47 0.35 0.58 0.23 0.45
Sub-MoE 0.25 0.58 0.58 0.46 0.25 - - -
Heapr 0.27 0.64 0.64 0.46 0.33 0.71 0.33 0.48
Qwen3-30B-A3B
0% Original 0.34 0.79 0.71 0.60 0.54 0.79 0.59 0.62
25% HC-SMoE 0.22 0.64 0.61 0.40 0.35 0.59 0.41 0.46
Sub-MoE 0.25 0.70 0.66 0.47 0.44 - - -
Heapr 0.33 0.77 0.70 0.55 0.49 0.78 0.50 0.59
50% HC-SMoE 0.13 0.44 0.50 0.29 0.23 0.44 0.32 0.34
Sub-MoE 0.23 0.68 0.63 0.41 0.40 - - -
Heapr 0.25 0.67 0.63 0.38 0.41 0.67 0.36 0.48
Qwen2-57B-A14B
0% Original 0.33 0.75 0.74 0.63 0.46 0.81 0.39 0.59
40% NAEE 0.31 0.73 0.73 0.55 0.46 0.76 0.36 0.55
MoE-I2 0.26 0.70 0.46 0.71 0.41 0.75 0.30 0.51
$D^{2}$-MoE 0.33 0.75 0.75 0.61 0.45 0.79 0.36 0.58
Heapr 0.33 0.75 0.74 0.64 0.46 0.81 0.39 0.59
4.1 Experimental Setup
Models and Setup.
We evaluate our approach on a broad spectrum of model architectures and scales to assess its generality and effectiveness, including DeepseekMoE-16B-Base (Dai et al., 2024), Qwen1.5-MoE-A2.7B-Chat (Team, 2024a), Qwen2-57B-A14B (Team, 2024b), and Qwen3-30B-A3B. All experiences are calibrated using Wikitext-2, with our method employing 128 sequences of 2048 tokens. Notably, our method introduces no additional tunable hyperparameters.
Baselines.
For our comparisons, we evaluate six recently proposed high-performance compression methods, including expert dropping (NAEE (Lu et al., 2024), MoE-I2 (Yang et al., 2024)), expert merging (MC-SMoE (Li et al., 2024), HC-SMoE (Chen et al., 2025)), and expert decomposition (Sub-MoE (Li et al., 2025a), $D^{2}$-MoE (Gu et al., 2025)). For fairness and transparency, the reported results in Table 1 are primarily drawn from the respective original papers, ensuring consistency.
Evaluation.
We report results on seven zero-shot benchmarks using the LM-Evaluation-Harness (version 0.4.7) (Gao et al., 2024), including HellaSwag (Zellers et al., 2019), Mathqa (Amini et al., 2019), OpenBookQA (OBQA) (Mihaylov et al., 2018), PIQA (Bisk et al., 2020), WinoGrande (Sakaguchi et al., 2021), ARC-Easy and ARC-Challenge (Boratko et al., 2018). These tasks collectively enable repeated and consistent evaluation of our method across varied domains and reasoning tasks.
4.2 Main Results
Compression Performance.
As shown in Table 1, Heapr achieves exceptional performance across various MoE models and compression ratios. Notably, our method delivers near-lossless compression. At pruning ratios of $20\%\sim 25\%$, Heapr matches the performance of the original models on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat. More impressively, on Qwen2-57B-A14B, Heapr maintains performance almost identical to the original model even at a high $40\%$ compression ratio. In contrast, competing new methods such as Sub-MoE, $D^{2}$-MoE, and NAEE show significant accuracy degradation under the same conditions. Furthermore, on the latest Qwen3-30B-A3B model, Heapr incurs only a minimal performance loss at a $25\%$ pruning ratio, with the average accuracy dropping slightly from $0.62$ to just $0.59$. These results strongly highlight the unique advantage of Heapr in pruning at the atomic expert level, enabling substantial model efficiency improvements while maintaining and preserving core model performance effectively.
Compare to CAMERA-P.
[…]
Performance Boundary of Pruning.
Figure 2: Performance of DeepSeekMoE-16B-Base under varying compression ratios, with corresponding FLOPs saving on WikiText2 data.
Figure 2 reports the performance of Heapr on DeepSeekMoE-16B-Base using a random 128-sample subset of WikiText-2 with $2048$ tokens under different compression ratios, where the ratio denotes the fraction of parameters removed relative to the full model size. For compression ratios below $0.4$, the pruned models retain over $93\%$ of baseline accuracy while already reducing FLOPs by $21\%$. In this regime, the accuracy curve remains nearly flat, revealing substantial redundancy among micro-experts and confirming that Heapr can effectively identify and remove them. As compression increases further, accuracy degrades gracefully, highlighting a clear trade-off between efficiency and performance. Even at an extreme compression ratio of $0.9$, the model preserves about $55\%$ of its baseline accuracy while achieving a $38\%$ reduction in FLOPs. These results demonstrate both the robustness of Heapr under moderate pruning and its effectiveness in enabling aggressive acceleration while retaining performance.
4.3 Ablations
Table 2: Comparison of layer-wise pruning (CAMERA-P and Heapr) versus global pruning (Heapr) on DeepSeekMoE-16B-Base and Qwen1.5-MoE-A2.7B-Chat across seven zero-shot tasks, with acc_norm reported for DeepSeekMoE-16B-Base and accuracy for others.
Ratio Method Openb. ARC_e WinoG. HellaS. ARC_c PIQA MathQA Average
DeepSeekMoE-16B-Base
20% CAMERA-P 44.00 71.80 70.17 75.02 45.56 78.62 31.46 59.52
Heapr-L 44.01 72.64 70.01 75.55 47.27 79.76 30.92 60.03
Heapr-G 44.80 73.73 71.43 76.57 47.01 79.82 31.42 60.68
40% CAMERA-P 43.20 70.71 68.51 69.04 42.24 75.41 29.01 56.87
Heapr-L 42.80 70.45 68.35 68.10 43.69 76.12 29.41 56.99
Heapr-G 41.40 72.05 69.06 70.79 45.05 76.39 29.85 57.80
Qwen1.5-MoE-A2.7B-Chat
25% Heapr-L 30.60 66.58 66.77 55.09 38.05 76.61 33.97 52,52
Heapr-G 31.80 68.60 67.22 55.67 37.56 76.39 34.87 53.59
50% Heapr-L 27.00 63.26 64.01 47.00 33.70 69.80 32.29 48.15
Heapr-G 27.01 63.89 64.32 46.35 34.22 70.86 33.37 48.57
Global vs. Layer-wise Pruning.
As shown in Table 2, layer-wise pruning (Camera-P, Heapr-L) ranks the importance of atomic experts within each MoE layer and prunes the bottom $r\%$, whereas global pruning (Heapr-G) ranks the importance of all atomic experts across the entire model. Compared with Camera-P, our layer-wise pruning Heapr-L achieves superior performance, indicating that the atomic expert importance metric, as derived from equation 13, provides a more effective pruning criterion within individual layers. Furthermore, Heapr-G, by leveraging global pruning and importance scores across all layers, achieves even stronger and more consistent results, validating the global consistency of the atomic expert importance thoroughly analyzed in Section 3.2.
Impact of Calibration Data.
Figure 3: Performance of DeepSeekMoE-16B-Base under a 20% compression ratio, using calibration data randomly sampled from WikiText-2 and C4.
Table 3 demonstrates that the performance of our Heapr algorithm is largely unaffected by the choice of calibration data, whether they are WikiText-2 or C4 dataset. This highlights the remarkable robustness and generalizability of our method, as it consistently performs well across different calibration corpora and domains. Furthermore, the table also explores the significant impact of calibration set size on pruning performance. As the number of calibration samples increases, the model’s performance improves consistently, indicating that larger calibration sets offer richer statistical coverage, which provides more reliable and informative signals for effective compression. These results strongly suggest that our method is not only robust to variations in calibration data but also benefits from the inclusion of additional diverse samples, further enhancing its overall effectiveness and stability.
5 conclusion
In this work, we introduce Heapr, a novel method that refines expert-level pruning in MoE models by enabling a more flexible and fine-grained pruning strategy at the atomic expert level. Inspired by the principles of the Optimal Brain Surgeon (OBS) theory, we evaluate the importance of atomic experts using second-order information. By transforming the analysis from the expert parameter space to that of atomic expert parameters, and further shifting it to the atomic expert output space, we significantly reduce the computational and storage bottlenecks associated with the second-order information matrix. Heapr requires only two forward passes and one backward pass to efficiently compute the importance of atomic experts. Extensive experiments on various modern MoE models demonstrate that Heapr outperforms state-of-the-art pruning methods, achieving near-lossless pruning with pruning rates of $20\%\sim 25\%$. More importantly, our method provides a much finer-grained perspective on MoE expert pruning, which we hope will contribute to a deeper, more comprehensive understanding of MoE models. Future work will explore large-scale experiments across a wider range of model and investigate the potential of parameter compensation methods after the pruning.
Appendix Overview
Appendix A Derivation of the importance of atomic expert
We provide a detailed derivation of the importance measure introduced in equation 13.
Consider the negative log-likelihood loss $\ell$, whose per-sample gradient with respect to the parameters $\Theta$ can be written as
$\nabla_{\Theta}\ell=\bm{J}_{k}^{\top}\mathbf{g}_{\mathbf{e}_{k}},$ (17)
where $\bm{J}_{k}\in\mathbb{R}^{d\times P}$ is the Jacobian of the atomic expert output $\mathbf{e}_{k}$ with respect to its parameters $\Theta$, and $\mathbf{g}_{\mathbf{e}_{k}}\in\mathbb{R}^{d}$ is the gradient of the loss with respect to $\mathbf{e}_{k}$. By definition, the Fisher Information Matrix is
$\bar{\bm{H}}=\bm{F}=\mathbb{E}\big[(\nabla_{\Theta}\ell)(\nabla_{\Theta}\ell)^{\top}\big].$ (18)
After model convergence, the Jacobian $\bm{J}_{k}$ can be treated as independent (Martens & Grosse, 2015), substituting the expression of $\nabla_{\Theta}\ell$ gives
$\bm{F}=\bm{J}_{k}^{\top}\,\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\bm{J}_{k}.$ (19)
Returning to the quadratic optimization problem in the OBS framework:
$\min_{\delta\Theta}\ \tfrac{1}{2}\,\delta\Theta^{\top}\bm{F}\,\delta\Theta\quad\text{s.t.}\quad\bm{J}_{k}\delta\Theta+\mathbf{e}_{k}=0,$ (20)
we define the auxiliary variable $\mathbf{u}=\bm{J}_{k}\delta\Theta$. The constraint becomes $\mathbf{u}+\mathbf{e}_{k}=0$, i.e., $\mathbf{u}=-\mathbf{e}_{k}$, and the objective reduces to
$\min_{\mathbf{u}}\ \tfrac{1}{2}\,\mathbf{u}^{\top}\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{u}\quad\text{s.t.}\quad\mathbf{u}+\mathbf{e}_{k}=0.$ (21)
Plugging in the constraint yields the optimal cost:
$\Delta\ell=\tfrac{1}{2}\,\mathbf{e}_{k}^{\top}\,\mathbb{E}\!\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{e}_{k}.$ (22)
We therefore define the importance of the $k$-th atomic expert as
$s\;=\;\tfrac{1}{2}\,\mathbf{e}_{k}^{\top}\,\mathbb{E}\left[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}\right]\mathbf{e}_{k}$ (23)
which is a scalar since $\mathbf{e}_{k}\in\mathbb{R}^{d}$ and $\mathbb{E}[\mathbf{g}_{\mathbf{e}_{k}}\mathbf{g}_{\mathbf{e}_{k}}^{\top}]\in\mathbb{R}^{d\times d}$. This formalizes equation 13 in the main text: the smaller the value of $s$, the less impact the $k$-th atomic expert has on the overall model loss, making it a better candidate for pruning.